OpenStudy (anonymous):

Solve the equation. Check the solution. -4/x+1=-1/x+5

OpenStudy (anonymous):

\[-4/x+1=-1/x+5\]

OpenStudy (anonymous):

-3/x = 4

OpenStudy (anonymous):

that's not one of my answer choices..

OpenStudy (uri):